# The kinematic cosmic dipole beyond Ellis and Baldwin

### v1

@@ 5 Conclusion
In this letter, we have demonstrated that the Ellis & Baldwin formula can be generalized to any class of light sources, regardless of their spectral shape or luminosity distribution, provided the coefficients $x$ and $\alpha$ are appropriately redefined. This result is essential for addressing the cosmic dipole anomaly, as it shows that sources other than quasars and radio galaxies can, in principle, be used to measure our velocity relative to the CRF.
We have also shown that, in the case of quasars, using spectra from AKARI, the resulting effective coefficient $\alpha$ within the CatWISE $W1$ band is generally consistent with the value obtained by fitting the spectrum with a pure power law. Moreover, we find that the procedure used in (Secrest et al. 2021), which relies only on the $W1$ and $W2$ magnitudes, may slightly underestimate the true spectral index by up to $\sim 0.16$. Still, this difference is not sufficient to significantly alter the conclusions of this work, that has shown that the quasar dipole is a factor of $\sim 2$ larger than the kinematic dipole expectation. Nevertheless, the limited number of available quasar spectra in this band prevents us from drawing definitive conclusions on this point. In the future, this issue may be clarified using spectroscopic libraries from missions such as SPHEREx (Crill et al. 2020).
In the general case, however, determining these new effective coefficients rigorously for a given dataset remains challenging, particularly for the effective spectral index $\alpha_{\rm eff}$. First, the spectral index is no longer independent of redshift. This is not necessarily problematic, as (von Hausegger 2024) showed that the original EB formula remains valid even when the indexes varies with redshift, $\alpha(z)$ and $x(z)$, provided some redefinitions. However, if this demonstration can be adapted to this general case of non-power laws is still an open question. Beyond this, another important issue is that most wide-area sky surveys are photometric, and the calculation of $\alpha_{\rm eff}$ requires a good quality of spectra within the range of the photometric band, and spectra are typically available for only a small subset of objects through cross-matching. This may introduce biases which will have to be carefully handled. Future applications of this test will need to address these issues directly.
@@ Acknowledgements.
The author thanks Roya Mohayaee, Nathan Secrest, Reza Ansari, Johann Cohen-Tanugi and Reiko Nakajima for their important insights and help.

### v2

@@ 5 Conclusion
In this letter, we have demonstrated that the Ellis & Baldwin formula can be generalized to any class of light sources, regardless of their spectral shape or luminosity distribution, provided the coefficients $x$ and $\alpha$ are appropriately redefined. This result is essential for addressing the cosmic dipole anomaly, as it shows that sources with more complex spectral profile than power laws such as galaxies in the near-infrared or visible light can in principle be used to measure our velocity relative to the CRF with photometric surveys. Upcoming large scale photometric surveys such as LSST (Ivezić et al. 2019) or Euclid (Euclid Collaboration et al. 2025) will provide us with the opportunity to apply the Ellis and Baldwin test to such data.
We have also shown that, in the specific case of quasars, using spectra from AKARI, the resulting effective coefficient $\alpha$ within the CatWISE $\rm W1$} band is generally consistent with the value obtained by fitting the spectrum with a pure power law. Moreover, we find that the procedure used in (Secrest et al. 2021), which relies only on the $\rm W1$} and $\rm W2$} magnitudes, may only slightly underestimate the true spectral index by up to $\sim 0.16$. Still, this difference is not sufficient to significantly alter the conclusions of this work, that has shown that the quasar dipole is a factor of $\sim 2$ larger than the kinematic dipole expectation. Nevertheless, the limited number of available quasar spectra in this band prevents us from drawing definitive conclusions on this point. In the future, this issue may be clarified using spectroscopic libraries from missions such as SPHEREx (Crill et al. 2020).
In the general case, however, determining this effective spectral index $\alpha_{\rm eff}$ rigorously for a given dataset remains challenging. First, the spectral index is no longer independent of redshift. Beyond this, another important issue is that most wide-area sky surveys are photometric, and the calculation of $\alpha_{\rm eff}$ requires a good quality of spectra within the range of the photometric band. Spectroscopic surveys have generally much less objects and this index may only be available for only a small subset of sources through cross-matching. Good quality spectra is generally only available for more luminous objects, therefore measuring $\alpha_{\rm eff}$ for sources at the flux limit specifically would be complicated. This may introduce biases which will have to be carefully handled. Future applications of this test will need to address these issues directly.
@@ Acknowledgements.
The author thanks Roya Mohayaee, Nathan Secrest, Reza Ansari, Johann Cohen-Tanugi, Reiko Nakajima and Sebastian Von Hausegger for their important insights and help.